OpenStudy (anonymous):

What is the domain and range of these functions y=-0.7x+1.9 and y=0.7x+1.9

OpenStudy (anonymous):

I don't get why the answers are: Left half: Domain: \[-2.7 \le x \le 0\] range: \[0 \le y \le 1.9\] Right half: Domain: \[0 \le x \le 2.7\] range: \[0 \le y \le 1.9\]